OpenStudy (anonymous):

f Indicate the equation of the given line in standard form. The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

OpenStudy (anonymous):

put the line in standart form first: y-4=2/3x-2 3y-12 =2x-6 2x-3y=6 then use vector [2,-3] to write new equation: 2(x-1)-3(y+2) =0 2x-2-3y -6=0 2x-3y=8

OpenStudy (anonymous):

thx man that one gave me alot of trouble aprreciate it! :)